A ball is dropped from a height of 126.8 meters. Its height above ground (in meters) t seconds later is given by: s(t)=−17t^2 +126.8


What is the instantaneous velocity of the ball when it strikes the ground?


1
Expert's answer
2021-12-16T15:41:03-0500
"v(t)=s'(t)=-34t"

"s(t)=0=>−17t^2 +126.8=0, t\geq0"

"t=\sqrt{126.8/17}"

"v=-34\sqrt{126.8/17}\approx-92.857\ m/s"

The instantaneous velocity of the ball when it strikes the ground is

"-92.857\ m/s" (downward speed with value of "92.857\ m/s" ).